OpenStudy (anonymous):

At the football game, the seniors sold slices of pizza for $.75each annd hamburgers for $1.35 each. They sold 40 more slices of pizza than hamburgers, and sales totaled $292.50. How many slices of pizza did they sell?

OpenStudy (curry):

1.35x+0.75(40+x)=292.50

OpenStudy (curry):

1.35+0.75x+30

OpenStudy (curry):

2.1x+30=292.50 262.50 =2.1x

OpenStudy (curry):

165 slices of pizza

OpenStudy (amistre64):

p = (292.50 - 40(1.35))/(.75-1.35)

OpenStudy (amistre64):

..... thats off, i read the question wrong

OpenStudy (curry):

am i right amis

OpenStudy (amistre64):

p = 40 + h p - h = 40 ; *y px + hy = z py - hy = 40y px + hy = z ------------- p(x+y) = z - 40y p = (z-40y)/(x+y) maybe? p = (292.50 - 40(1.35))/(.75+1.35) maybe :)

OpenStudy (amistre64):

well that dint turn out right lol

OpenStudy (curry):

my way was easier amis

OpenStudy (curry):

look at mty work

OpenStudy (anonymous):

yes, i agree curry

OpenStudy (anonymous):

okay...Tell me if i got this right, it should be easy. What is the lenght of the diagonal of a rectangle with width 8 cm and length 6 cm? My answer: A=lw 8x6=48

OpenStudy (curry):

its 10 not 48

OpenStudy (anonymous):

Amis language is too complicated for me...-_-

OpenStudy (amistre64):

p-40 = h p(.75) + (p-40)(1.35) = 292.50 p(.75) + p(1.35) - 40(1.35) = 292.50 ... got it :) p = (292.50 +40(1.35))/(.75+1.35) = 165 :) i had a formulae for something similar and had to tweek it for my own personal satisfatcion

OpenStudy (amistre64):

why aint this formating right?

OpenStudy (amistre64):

test test test these are on different lines

OpenStudy (curry):

nvm

OpenStudy (amistre64):

its +40(1.35) .. i had to do it the long way to see where my adjustments were :)